the room is 30 feet long 24 feet wide, and 9 feet high. Find the area and the volume of the room

Answer by nerdybill(7384)

the room is 30 feet long 24 feet wide, and 9 feet high. Find the area and the volume of the room
.
Area = width * length = 24*30 = 720 square feet
.
Volume = width * length * height = 24*30*9 = 6480 cubic feet